At the end of the year, a company offered to buy 5,000 units of a product from X Company for a special price of $11.00 each instead of the company's regular price of $19.00 each. The following information relates to the 65,000 units of the product that X Company made and sold to its regular customers during the year:

Per-Unit Total     
Cost of goods sold $9.63    $625,950   
Period costs 2.54    165,100   
Total $12.17    $791,050   


Fixed cost of goods sold for the year were $139,100, and fixed period costs were $74,750. Variable period costs include selling commissions equal to 3% of revenue.

6. Profit on the special order is

Incorrect. Tries 1/3 Previous Tries


7. Assume the following two changes for the special order: 1) variable cost of goods sold will increase by $0.86 per unit, and 2) there will be no selling commissions. What would be the effect of these two changes on the special order profit?

Tries 0/3


8. There is concern that regular customers will find out about the special order, and X Company's regular sales will fall by 850 units. As a result of these lost sales, X Company's profits would fall by
Answer #1
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Amount
Revenue from sales = 5000*11            55,000.00
Less costs
Variable cost of goods sold = (625950 - 139100)/65000*5000            37,450.00
Variable period costs = (53,300)/65000*5000 + 55000*3%              5,750.00
Total costs            43,200.00
Profit from special order            11,800.00
7)
Revenue from sales = 5000*11            55,000.00
Less costs
Variable cost of goods sold = 37,450 +5000*.86            41,750.00
Variable period costs = (53,300)/65000*5000              4,100.00
Total costs            45,850.00
Profit from special order              9,150.00
Change in Profit = 11,800 - 9150              2,650.00
8)
Revenue from 850 units = 850*19            16,150.00
Less COGS = 850*7.49              6,366.50
Less period cost = 850*1.39              1,181.50
X Company's profits would fall by              8,602.00
Variable cost of goods sold per unit= (625,950-139,100)/65000                      7.49
Total period costs         165,100.00
Fixed period costs            74,750.00
Variable period costs            90,350.00
Variable period costs per unit =90,350/65000                      1.39
Selling commission =65000*19*3%            37,050.00
Other variable costs = 90350 - 37050            53,300.00